### v1

# pH-Driven Fluorescence in Solid-State Lysine

Abstract
There is currently a growing interest in understanding the origins of intrinsic fluorescence as a way to design non-invasive probes for biophysical processes. In this regard, understanding how pH influences fluorescence in non-aromatic biomolecular assemblies is key to controlling their optical properties in realistic cellular conditions. Here, we combine experiments and theory to investigate the pH-dependent emission of solid-state L-Lysine (Lys). Lys aggregates prepared at different pH values using HCl and H2SO4 exhibit protonation- and counterion-dependent morphology and fluorescence, as shown by microscopy and steady-state measurements. We find an enhancement in the fluorescence moving from acidic to basic conditions. To uncover the molecular origin of these trends, we performed non-adiabatic molecular dynamics simulations on three Lys crystal models representing distinct protonation states. Our simulations indicate that enhanced protonation under acidic conditions facilitates non-radiative decay via proton transfer, whereas basic conditions favor radiative decay. These findings highlight pH as a key factor tuning fluorescence in Lys assemblies, offering insights for designing pH responsive optical materials based on non-aromatic amino acids.
keywords:
[…]
Fluorescence in biological materials is typically attributed to aromatic amino acids due to their conjugated -systems and favorable photophysical properties1. However, recent findings have revealed intrinsic photoluminescence in non-aromatic amino acids and peptides, particularly under aggregation in solution or in the solid state.2, 3, 4, 5, 6, 7, 8 A key mechanism proposed to underlie this phenomenon is aggregation-induced emission (AIE),9 wherein non-emissive species in dilute solution become strongly emissive upon aggregation. While AIE is well established in synthetic organic luminogens,10, 11, 12 its manifestation in non-aromatic biomolecular assemblies, such as peptide and amino acid aggregates, has opened new directions for exploring photophysics in biologically relevant systems.13, 14, 15, 16
Hydrogen bonding plays a central role in mediating luminescence in non-aromatic systems, influencing both the extent of electron delocalization and which vibrational modes lead to non-radiative decay on the excited state17, 2, 18, 6. Environmental factors such as pH, solvent polarity, and ionic strength can profoundly affect these interactions, often altering morphology and photophysical properties19, 20. However, how these factors reshape the excited state (ES) potential energy landscape remains unclear21. While peptide self-assembly under neutral pH has been extensively studied,22, 23, 24 less attention has been given to extreme pH conditions, which can drastically alter conformation, charge distribution, and intermolecular interactions25, 21. In this regard, pH is known to play an important role in tuning fluorescence in aromatic or more generally conjugated molecular systems26, 27, 28, 29. Thus, the potential of designing novel probes based on intrinsic fluorescence as indicators of pH in optical sensing and imaging, could open up new applications in biological and material sciences29, 30.
In this context, L-Lysine (Lys), a non-aromatic amino acid with a protonatable side chain, provides a compelling model to probe pH-dependent luminescence. Depending on pH, Lys adopts distinct protonation states (see Fig. 1),31 which in turn strongly influence its self-assembly and crystal packing, thereby modulating its photophysical properties.32, 4. Similar effects have been observed in poly-lysine systems, where pH influences aggregate morphology and optical response.33, 34, 35. Recent work by Stagi et al.32 combined experiments and theory to investigate the photophysical properties of self-assembled Lys in aqueous solution. Absorption spectra showed a distinct band at 278 nm, attributed to electronic transitions in hydrogen bonded Lys clusters. Classical molecular dynamics (MD) simulations further confirmed that Lys forms clusters of varied size through hydrogen bonds (HBs) and hydrophobic interactions, with the former playing a dominant role. Fluorescence measurements in solution exhibit strong pH dependence: emission is enhanced under neutral and basic conditions but quenched at low pH, where high protonation is proposed to inhibit functional group interactions, reducing emission efficiency.32 However, the precise ES relaxation pathways that affect fluorescence emission in titratable amino acids, such as lysine, remains poorly understood.
In this contribution, by combining both experiments and theory, we address a critical gap in our understanding of how pH affects non-aromatic fluorescence using solid-state lysine as a model system. Solid-state Lys aggregates were prepared at four pH values (1, 4, 7, and 10) to probe how the protonation influences both morphological and photoluminescence properties. Since pH was adjusted with either a monoprotic (HCl) or a diprotic (H2SO4) acid, we also assessed the impact of the counterion on these properties. Morphology was characterized by optical and scanning electron microscopy (SEM), while photoluminescence was studied via confocal microscopy, steady-state fluorescence, and external quantum efficiency (EQE) measurements. Our results show that both protonation state and the selected acid significantly influence aggregate morphology and fluorescence, establishing these as key parameters for tuning optical properties. To uncover the molecular origins behind these trends, we performed non-adiabatic molecular dynamics (NAMD) simulations to follow the ES dynamics of Lys. While direct simulation of the experimental aggregates is precluded by the lack of precise structural data, we can instead model crystal structures of Lys representing distinct protonation states (Lys2+, Lys+, and Lys0 in Fig. 1).36, 37, 38 These models offer well-defined frameworks to interpret fluorescence behavior and examine ES processes. Our simulations indicate that fluorescence decreases from low to high protonation states (), consistent with our experimental data and prior reports32, 4. Mechanistically, increased protonation promotes proton transfer (PT) processes that enable non-radiative decay, causing fluorescence quenching under acidic conditions.
As mentioned above, we investigated the morphology of solid-state Lys aggregates formed at pH 1, 4, 7, and 10, adjusted using either HCl or H2SO4, by optical and scanning electron microscopy. The characterization was performed on Lys samples prepared in water at the concentration of g/mL ( mol/L) corresponding to the maximum of its solubility. This concentration–which is approximately tenfold higher than the critical aggregation concentration (CAC)–ensures the formation of aggregated clusters (see Fig. S1 and discussion in the Supporting Information (SI)).
Fig. 2 shows clear morphological differences depending on both acid type and pH. Under strongly acidic conditions (pH = 1), both HCl- and H2SO4-treated samples initially formed oily films at room temperature, with no macroscopically ordered structures observed over time. However, upon overnight cooling to , well-ordered crystals emerged (Fig. 2a, Lys + HCl; Fig. 2e, Lys + H2SO4). Optical microscopy revealed distinct morphologies: HCl-treated samples formed smaller, densely packed crystals, likely due to enhanced nucleation promoted by chloride ions. In contrast, H2SO4 promoted the formation of larger, elongated crystals, suggesting that sulfate anions favor fewer nucleation events, leading to more sparsely distributed structures. Reheating caused crystals to revert to oily films, preventing further morphological study.
At pH 4, HCl-treated Lys aggregates (Fig. 2b) formed thin and disperse small fibers, whereas H2SO4-treated aggregates (Fig. 2f) resulted in denser fibrous networks with a more compact appearance. This contrast suggests that sulfate ions can promote stronger cross-linking, leading to a more interconnected aggregation. At pH 7, HCl-derived Lys aggregates (Fig. 2c) formed a layered, sheet-like morphology, while those acidified with H2SO4 (Fig. 2g) a more porous, interconnected network, suggesting that sulfate anions promote a web-like assembly rather than compact layers. At pH 10, both acids yielded fibrous architectures (Fig. 2d, and Fig. 2h), with thinner and more interconnected fibers in HCl-treated samples. Cross-polarized light microscopy (Fig. S2) revealed birefringence, indicating isotropic organization of the fibrillar network at this pH.
Overall, optical and scanning electron microscopy reveal that both pH and acid type influence Lys aggregate morphology. To investigate how these structural differences impact optical properties, we turned to fluorescence-based techniques, starting with confocal and fluorescence microscopy, followed by steady-state spectroscopy and quantum efficiency measurements to better quantify the pH dependent fluorescence emission.
Confocal microscopy of Lys aggregates prepared with HCl (Fig. S3) and H2SO4 (Fig. 3) confirmed the evolution from crystalline‑like structures at pH 1 to diffuse fibrillar networks at pH 4, 7, and 10. This trend is evident in the far-left subpanels of Fig. 3b–d (dark-field images). For pH 1 (Fig 3a), optical microscopy is shown due to the absence of dark-field data. Each panel also includes fluorescence microscopy images recorded in the blue (second column), green (third column), and red (rightmost column) channels. A striking enhancement in fluorescence intensity is observed with increasing pH for both acids: aggregates at pH 1 exhibit weak or negligible emission across all channels, while those at neutral and basic conditions (Fig. 3b–d) show stronger fluorescence.
To quantify this, we collected steady-state fluorescence spectra across excitation wavelengths from 330-410 nm (10 nm step). The resulting spectra for Lys acidified with HCl and H2SO4 are shown in panels a-d and panels e-h of Fig. S4, respectively. Interestingly, we observed a linear correlation between excitation and emission wavelengths (Fig. S5 of the SI), consistent with previous findings in Lys aggregates in solution32, and poly-lysine polymers33, where this dependency was attributed to the ability of Lys to form aggregates of various size. Excitation-dependent fluorescence, or red-edge excitation shifts, has been observed in a variety of systems, including carbon nanodots39, sodium-uracil thin films40, self-assembled peptide nucleic acids41, and organic fluorescent nanoparticles42. Proposed mechanisms include heterogeneous aggregate populations39, slow environment relaxation around the excited fluorophore41, 40, and collective proton motions within HB networks that modulate the ES manifold42. These findings suggest that the excitation-dependent emission in our Lys samples likely arises from similarly heterogeneous supramolecular environments, shaped by protonation state and ionic interactions.
To isolate the role of protonation and acid identity, we compared normalized fluorescence spectra at nm for all samples, as shown in Fig. 3e-f. Both acid treatments yielded broad emissions (350–550 nm), with emission maxima () centered between 400 nm and 420 nm, depending on pH and acid. For HCl (Fig. 3e), blue-shifts from 422 nm (pH 1, red line) to 403 nm (pH 10, cyan line), while for H2SO4 (Fig. 3f), the pH 1 sample shows the most blue-shifted emission ( nm), and all other pH values exhibit nearly identical spectra ( 414 nm). These subtle spectral differences align with SEM observations: H2SO4 yields more consistent fibrous morphologies across pH, while HCl shows greater variation, potentially explaining the progressive red-shift.
While these spectra offer an initial view of how pH influences fluorescence, their normalization masks variations in absolute emission intensity. For this reason, we measured external quantum efficiency (Fig. 3g), since it accounts for light extraction efficiency43 and is therefore particularly relevant for assessing emission in solid-state devices44, 45, 46. Here, a striking pH dependence emerges: EQE increases from at pH 1 for both acid treatments and rises to nearly 5 at pH 7 for Lys with HCl (green plot) and at pH 7-10 for Lys with H2SO4 (purple plot), reflecting higher emission efficiency under low protonation. These trends are consistent with previous studies on solution-phase Lys aggregates, which linked protonation to fluorescence quenching32. Notably, the two acids impart distinct efficiencies: with H2SO4, EQE increases from pH 1 to pH 7 and then plateaus at pH 10, whereas with HCl the efficiency peaks at pH 7 and declines at pH 10. This divergence underscores the additional role of the counterion, beyond protonation alone, in dictating aggregate structure and, consequently, solid‑state fluorescence efficiency.
Our experimental characterization clearly shows that both pH and counterion modulate the emission of solid‑state Lys aggregates, highlighting the influence of protonation and aggregation on fluorescence. However, these techniques cannot directly resolve the molecular origins of this effect, specifically, how protonation state shapes the electronic structure and ES dynamics. To gain molecular insight into this, we now move to discussing the theoretical results that emerge from our non-adiabatic molecular dynamics simulations performed on three Lys crystal models corresponding to distinct protonation states, as shown in Fig. 4.
Structural differences among the lysine crystals reflect protonation-dependent variations that mirror the pH effects observed in aggregate morphologies, supporting their use as proxies to study solid-state Lys photophysics. As we will see shortly, these models also correctly reproduce the trends in the emission observed as a function of pH. L-Lys sulfate (LLS, Fig. 4a) adopts a folded conformation stabilized by a dense HB network between SO and the -NH and CO2H groups of lysine36, 47, with HBs exclusively between Lys and sulfate and no direct Lys-Lys interactions due to electrostatic repulsion. In L-Lys monohydrochloride dihydrate (LLMHCl, Fig. 4b), all amines are protonated, the C-terminus is deprotonated, and both Lys–Lys and Lys–water HBs are observed. Cl- counterions further stabilize the structure via electrostatic interactions with both the - and -amino groups37, 48. The zwitterionic L-Lys hemihydrate (LLH, Fig. 4c) forms layers of head-to-head hydrogen bonded dimers through -NHCO interactions, a network of interactions previously linked to intrinsic fluorescence38, 49, 4.
For each crystal, we performed 500 fs NAMD simulations, within a QM/MM framework, on 50 statistically independent initial conditions (see the Computational Details section in the SI for details). In Fig. 4a–c, the lysine dimers (rendered with glossy CPK spheres) represent the QM regions used in our simulations. After 500 fs, most trajectories remained in the first ES for LLH (pH = 10) and LLMHCl (pH = 7), namely 98 and 92, respectively, whereas only 28 remained in S1 for LLS (pH = 1), as shown in Fig. 4d (orange plot). This severe quenching at low pH mirrors the experimental EQE trend observed in Fig. 3f, reproduced here for direct comparison, highlighting the key role of protonation in modulating the ES lifetimes.
In Fig. 4e, we report the fluorescence spectra of LLS (red line), LLMHCl (green line), and LLH (cyan line) averaged over the trajectories remaining in S1. Vertical lines indicate the averaged lowest absorption (excitation) transitions: blue (LLS), magenta (LLMHCl), and yellow (LLH). For all systems, excitation wavelengths cluster near nm, exhibiting a 120 nm ( eV) blue-shift relative to experiment. This discrepancy likely arises from a combination of known TDDFT limitations, such as the tendency of CAM-B3LYP to overestimate vertical excitation energies50, as well as neglect of nuclear quantum effects (NQEs), and structural defects or disorder51, 52. Several previous studies have shown that the inclusion of NQEs can alter both the peak absorption in organic systems by up to 1 eV, as well as affect the line-shape53, 54, 2, 55.
Across protonation states, the transition consistently occurs from the highest occupied molecular orbital (HOMO) to the lowest unoccupied molecular orbital (LUMO), which exhibit n and character, respectively. An example for LLS is shown in Fig. 4f: the n-type HOMO localizes on the C-terminal carboxyl and partly on SO oxygen atoms, while the -type LUMO is on the same Lys monomer, primarily on the C-terminal oxygen and along the CCα bond. Similar localizations occur for LLMHCl (Fig. S6a) and LLH (Fig. S7a), with the n-type HOMO on CO oxygens and the -type LUMO on this group and adjacent Cα. Cl- ions show no active role in LLMHCl. The electron density alternates between the two QM Lys monomers, consistent with dimer symmetry, and localizes within H-bonding regions, linking local non-covalent interactions with photoexcited electronic properties.
While the excitation energies show little pH dependence, emission spectra reveal more pronounced differences. The computed emission trend in Fig. 4e mirrors the one observed experimentally for Lys samples acidified with H2SO4 (Fig. 3f). The crystal representing pH 1 (LLS) shows the most blue-shifted emission ( nm), followed by LLMHCl (pH 7) and LLH (pH 10), both exhibiting emission maxima near 356 nm. Notably, the emission spectrum of LLMHCl displays an additional broad feature between 460–480 nm, originating from structures in which excited state distortions–particularly CCO2 deplanarization–are partially activated but insufficient to reach the conical intersection (CoIn). This mode, typically associated with non-radiative decay6, in this case gives rise to red-shifted emission from distorted yet emissive geometries. LLS emission instead exhibits a shoulder near the excitation energy, suggesting an ES minimum near the Franck-Condon region. Although emission energies remain overall blue-shifted by 50–80 nm (0.5-0.7 eV) relative to experiment, this discrepancy is significantly reduced compared to excitation, and spectral tails at longer wavelengths ( nm) are well captured for LLMHCl and LLH. In all three systems, the electronic transition retains the same n- character and orbitals localization as in the GS excitation (see Figs. S6-S8 in the SI).
Since the calculated emission spectra in Fig. 4e are normalized, this does not directly allow for a comparison of the intensities. We therefore investigated the oscillator strength (f) distributions for further insight. Normalized histograms and corresponding Gaussian fits for LLS, LLMHCl, and LLH are shown in Fig. S9a-c, with an overlay of the fitted curves in Fig. S9d. For the more fluorescent LLH (cyan) and LLMHCl (green), distributions peak below 0.001, with minor tails up to 0.005. These values indicate modest fluorescence intensity, well below that of typical aromatic fluorophores56, yet sufficient for detection (Fig. 4d). In contrast, LLS (red) shows a broader distribution with oscillator strengths reaching up to 0.01; however, the fluorescence signal is minimal due to the predominance of non-radiative decay pathways.
Non-radiative decay on the excited state essentially involves fluctuations or distortions along vibrational modes that lead to the CoIn and subsequently, decay to the ground state. Previous studies2, 57, 3, 18, 6 have shown that proton transfer (PT), carbonyl stretching, and deplanarization of the amide group act as critical modes in modulating the ES lifetimes of peptide chains, amyloid-like peptides, and glutamine crystals. Constraining these modes raises the energy barrier to the conical intersection, prolonging the ES lifetime. More recently, Banerjee and co-workers7 linked the weaker fluorescence of cysteine crystals grown in H2O, as opposed to D2O, to activation of CαN and SH stretching modes. Guided by the n character of LLS (Fig. 4f), we first focused on vibrational modes involving the C-terminus, as shown in Figs. S10-S11 for the two QM Lys monomers. We then investigated the coordinates associated with the various HBs present in the system (Fig. S12 in the SI). From this combined set of vibrational modes, a small number of coordinates, shown in Fig. 5, emerge as the dominant funnels to the CoIn.
The collective motion driving non-radiative decay in LLS (Fig. 5a) can be decomposed into two main contributions: (i) stretching of the CO bond coupled with deplanarization of the CαCO2 group, and (ii) PT coordinates involving the NH or OH groups of Lys and the sulfate oxygen. As shown in Fig. 5b, the CO bond elongates in S1 (orange distribution), increases further at the CoIn (dashed line), and contracts upon relaxation to S0 (blue distribution). A similar trend is observed for the second CO bond of the carboxylate group (Fig. S11c). This distortion is concerted with deplanarization of the amide group in S1, followed by replanarization after decay (Fig. S11d). These features are consistent with previously reported ES dynamics in non-emissive amino acid crystals, such as L-glutamine6.
However, unlike glutamine–where CO stretching is triggered by transient strengthening (PT coordinate ) of a -NHCO hydrogen bond between monomers–LLS exhibits HBs exclusively with sulfate counterions. As a result, CO bond activation and proton transfer do not necessarily occur within the same structural unit. The primary HB-related mode promoting non-radiative decay involves transient weakening of an -NHSO HB in S1, particularly near the CoIn, with PT typically occurring upon relaxation to S0 (Figs. 5c and S12f). A secondary PT mode, shown in Fig. 5d, involves the C-terminal OH group and displays a slight transient HB strengthening in S1, with occasional proton sharing or transfer to the sulfate oxygen at the CoIn, and reversion after decay. A significant number of trajectories that reach the CoIn show PT coordinates below Å, consistent with a weakening of the interaction. Collectively, these results highlight that CO bond distortions and HB fluctuations actively govern the accessibility of the CoIn under acidic conditions, with the sulfate anion also playing a crucial role in this dynamic behavior.
Similar CO elongation and CαCO2 deplanarization appear in the rare non-radiative trajectories of LLMHCl and LLH (Figs. S13–S14b–d), but no PT occurs in S or S. In both systems, weakening of the NHO coordinates is observed, with values often exceeding Å, but stems from other structural distortions driving the non-radiative decay. In LLMHCl, CO stretching and deplanarization lead to CαC elongation in S (Fig. S13e), distancing the acceptor from the NH donor (Fig. S13f). In LLH, significant CαN stretching near the CoIn (Fig. S14f) weakens one HB, while the other is disrupted by flipping of the carboxylate group (Figs. S14g–h). This CαN activation was recently implicated in the weak fluorescence of cysteine crystals.7 These decay events, although rare, highlight that under neutral and basic conditions, non-radiative decay is primarily driven by backbone distortions rather than HB dynamics. In contrast, under acidic conditions, the combination of enhanced protonation and the presence of flexible, multiply charged counterions–such as sulfate ions–increases the dynamical complexity of the hydrogen bonding network, opening new non-radiative pathways.
In conclusion, we have combined experiments and theory to investigate the evolution of fluorescence in solid-state lysine as a function of pH. To the best of our knowledge, the physical factors and mechanisms controlling intrinsic fluorescence as a function of pH in non-aromatic supramolecular assemblies remain poorly understood and have been explored in only a handful of studies2, 32, 21. One of our key findings is that both pH and the counterions introduced by the acids used to adjust it modulate protonation states and hydrogen bonding networks, altering the physical and chemical properties of the assemblies.
Using a synergy of microscopy techniques, we reveal subtle yet distinct morphological changes in lysine aggregates from acidic to basic conditions, which correlate with enhanced fluorescence intensity at higher pH. To rationalize these trends, we performed excited state molecular dynamics simulations on lysine crystal models, revealing that non-radiative decay pathways are more accessible at lower pH, consistent with experiments. Specifically, under acidic conditions, the presence of a counterion forming strong hydrogen bonds with the amino-acid, activates channels–including proton transfer–that facilitate easy access to the conical intersection.
Our findings open up important perspectives and challenges on both experimental and theoretical fronts. Specifically, advancing experimental methods to better characterize local and macroscopic structures of titratable amino acid aggregates as a function of pH remains a crucial factor to guide computational modeling. While our models capture qualitative trends, a unique aspect we uncovered in this study is that the non-radiative decay paths under acidic conditions involve proton transfer processes for both normal and strong hydrogen bonds. An important implication of this is that the choice of the QM/MM boundary may suppress or enhance proton transfer events. While we have found that our conclusions are robust to this factor (see Fig S15 in the SI), efforts that couple reactive QM/MM boundaries for excited state dynamics will be interesting to pursue in the future.
MM, GDM, DB, and AH acknowledge the European Commission for funding on the ERC Grant HyBOP 101043272. They also thank CINECA supercomputing (project NAFAA-HP10B4ZBB2) and MareNostrum5 (project EHPC-EXT-2023E01-029) for computational resources. CD, AA, and LV acknowledge the project PRIN2022TSLMHR entitled “Biomaterials from peptide self-assembling generated by mimicking protein amyloid-like structures”, funded by Ministero dell’Università e della Ricerca - NextGenerationEU (European Union). LC acknowledges Ricerca Corrente Project.
Conflict of Interest
The authors declare no conflict of interest.
Data Availability Statement
Additional experimental and computational details; determination of the critical aggregation concentration; birefringence microscopy images of Lys aggregates acidified with HCl and H2SO4 at pH 10; confocal fluorescence images of Lys aggregates prepared using HCl; normalized steady-state fluorescence spectra of Lys aggregates under varying excitation wavelengths (HCl and H2SO4); excitation-emission correlation plots; molecular orbitals involved in excitation and emission of L-Lysine sulfate (LLS), L-Lysine monohydrochloride dihydrate (LLMHCl), and L-Lysine hemihydrate (LLH); oscillator strength distributions for the three Lys crystal forms; vibrational coordinate analyses from trajectories undergoing non-radiative decay (LLS, LLMHCl, LLH); validation of the QM/MM setup for LLS; and additional references.
S1 Experimental Details
S1.1 Chemicals
L-Lysine (Lys) monohydrate, HCl (wt) and H2SO4 (wt) were purchased by Thermo Fisher Scientific (Milan, Italy).
S1.2 Samples preparation
All solutions were prepared by weight dissolving Lys in doubly distilled water; their pH was adjusted using HCl (wt) and H2SO4 (wt). pH was measured by pH-meter (Mettler Toledo FiveEasy Plus pH meter FP20, equipped with a micro electrode).
S1.3 Fluorescence spectroscopy
Fluorescence spectra were collected at room temperature on a Jasco spectrofluorophotometer (Model FP-8550), equipped with a continuous output Xenon arc lamp. Measurements in solution were conducted in a quartz cell with a 1.0 cm path length. Solid state fluorescence and photoluminescence quantum yields (PLQYs) were evaluated using an ILF-835 integrating sphere. PLQY was measured by the absolute method thereby overcoming the issue of unavoidable differences in the relative emission intensities arising from poorly reproducible drop-casted amino acid clusters. Fluorescence spectra of Lys at the solid state were acquired in a window up to 600 nm, probing the 330–410 nm excitation range at 10 nm steps. All acquisitions were averaged over five measurements after spectral correction.
S1.4 Confocal microscopy
After drop-casting on a microscope slide the Lys water solutions at various states of protonation and letting them dry overnight, samples were analyzed using a and magnifications on the MICA microhub (Leica, Wetzlar, Germany) platform. Fluorescence images were obtained using filters in the spectral regions of blue (, ), green (, ), and red (, ).
S1.5 Optical microscopy
To prepare the samples, 20 L of each Lys solution was placed onto a glass microscope slide and air-dried overnight at room temperature. The obtained samples were examined using an Optech BM80Pol microscope (Milan, Italy) for both cross polarized and bright-field light.
S1.6 Scanning Electron Microscopy
Morphological analysis of samples was carried out by field emission scanning electron microscope (ThermoFisher Scientific Phenom Pro G6). 10 L of Lys solutions were drop-casted on an aluminum stub and air-dried. A 7 nm coat of gold was sputtered at a current of 25 mA for 75 sec. The gold coated samples were then introduced into the specimen chamber and the images were acquired in high vacuum mode, at an accelerating voltage of 15 kV, through the Secondary Electron Detector (SED).
S1.7 Critical Aggregation Concentration Measuraments
Critical aggregation concentration (CAC) value for Lys in water was determined by fluorescence spectroscopy. In detail, two different fluorophores: a) 8-anilino-1-naphthalene sulfonic acid ammonium salt (ANS) and b) pyrene (pyr) were separately titrated with increasing amounts of the amino acid water solution. Fluorescence spectra were collected at room temperature on a Jasco spectrofluorophotometer (Model FP-8550), acquiring measurements in a quartz cell with a 1.0 cm path length at room temperature. The other experimental settings are: excitation and emission bandwidths = 5 nm; recording speed = 120 nmmin-1, excitation wavelength = 350 nm and 335 nm for ANS and pyr, respectively. The measurement was performed by adding small aliquots of the amino acid in 200 L of 200 M ANS or 10 M pyr water solution. At the end of the titration, the blank was subtracted. Fluorescence spectra were corrected for the blank and adjusted for the dilution. Tendency lines were extracted using a least squares method extrapolation. Titrations were conducted in triplicate. In the ANS method, the CAC value was determined by linear least-squares fitting the fluorescence emission at 475 nm, upon excitation at 350 nm versus the amino acid concentration as previously reported58. In the pyr method, the CAC value was determined by the break-point of the fitting of I1/I3 upon excitation at 335 nm versus the amino acid concentration59.
S2 Computational Details
S2.1 Classical molecular dynamics
Lys systems corresponding to different protonation states were modeled using crystallographic structures of L-Lysine sulfate (LLS), L-Lysine monohydrochloride (LLMHCl), and L-Lysine hemihydrate (LLH). Supercells of appropriate size were built for each system: 1588 for LLS, 11511 for LLMHCl, and 7116 for LLH. Classical molecular dynamics (MD) simulations were performed using the AMBER MD package version 202360, 61 with the Generalized Amber Force Field (GAFF) 62 applied to all molecular mechanics (MM) atoms. Each system was first equilibrated in the NVT ensemble at 300 K for 10 ns using a 2 fs time step and a Langevin thermostat63 with a collision frequency, , of ps-1. Next, a 50 ns production run was performed under the same conditions, with all bonds involving hydrogen atoms constrained. From this trajectory, 50 independent frames were extracted at 1 ns intervals and used as starting geometries for the subsequent quantum mechanics/molecular mechanics (QM/MM) simulations.
S2.2 Ground state QM/MM
To accurately capture the luminescent properties of the Lys crystals, it is essential to account for the influence of the surrounding environment. As fully QM treatment of the entire system is computationally prohibitive, we employed a hybrid QM/MM approach64, 65. For each crystal, we selected relevant QM regions (dimers; see Fig. 4a–c in the Main Text) from the 50 uncorrelated frames extracted from classical MD simulations. Thermal equilibration of each QM region at 300 K was performed using ORCA 5.0.466, 67, 68, 69, 70 (QM) interfaced with AMBER 202360, 61 (MM). The QM region was treated using the CAM-B3LYP exchange-correlation (xc) functional71 and the def2-SVP basis set72, 73 with the RIJ-COSX approximation74 to accelerate Coulomb and exact exchange integral evaluations. Equilibration was carried out for 1 ps in the NVT ensemble at 300 K using a Langevin thermostat63 with tight coupling ( ps-1) and no bond constraints. A 0.5 fs time step was used for 1 ps, followed by an additional 1 ps with a standard thermostat coupling ( ps-1). Final nuclear coordinates and velocities from these equilibrated QM/MM trajectories were used as initial conditions for the non-adiabatic molecular dynamics (NAMD) simulations described below.
In addition, vertical excitation energies were computed from these equilibrated structures to characterize the absorption spectra and identify relevant transitions. These time-dependent DFT (TDDFT) calculations were performed using the same ORCA/AMBER QM/MM framework, but with no MD propagation (0 MD steps), and the same xc functional, basis set, and approximations described above. For each Lys crystal system, the ten lowest singlet excited states were computed for 50 equilibrated initial conditions. Excitation plots shown in Fig. 4e in the Main Text correspond to the averaged energies of the lowest-energy absorption transition across these 50 frames. Molecular orbitals involved in these transitions were obtained using a post-processing tool available in ORCA.
S2.3 Non-adiabatic molecular dynamics
NAMD simulations were performed using the Trajectory Surface Hopping (TSH) method, specifically employing the SHARC formalism75, 76, which extends Tully’s Fewest Switches Surface Hopping (FSSH)77 to incorporate both time-dependent non-adiabatic couplings (TD-NACs) and spin–orbit couplings on an equal footing. In this approach, nuclei evolve classically on quantum potential energy surfaces computed on-the-fly. For further background, we refer the reader to comprehensive reviews of TSH methodology.78, 79, 80, 81.
All simulations–50 for each system–included the first 5 singlet states () to adequately describe the lowest excited states relevant for fluorescence studies. Each trajectory was initialized on the first excited state () and propagated in the NVE ensemble for 500 fs with a time-step of 0.5 fs using the velocity Verlet integration algorithm82, 83. The functional and basis set used in the ground-state QM/MM dynamics were kept unchanged in the NAMD simulations. Excitation energies and oscillator strengths were calculated at the TDDFT84 level, within the Tamm-Dancoff approximation (TDA)85, which improves TDDFT behavior near conical intersections86, 87, 88. NACs were computed via the Hammes-Schiffer–Tully scheme89 based on wavefunction overlap integrals between successive time steps. Electronic dynamics were integrated using 25 sub-steps per nuclear time step, and decoherence effects were included using the energy-based correction of Granucci et al.90 with a 0.1 Hartree parameter.
S2.4 Validation of the QM/MM scheme
The definition of the QM region is a critical factor for reliably capturing excited-state (ES) dynamics, particularly for proton transfer (PT) processes, which are known to be sensitive to the QM/MM boundary in flexible or solvated systems.96, 97 Although we focus on crystals, where structural rigidity simplifies this choice6, we assessed the robustness of our conclusions by expanding the QM region in the LLS system to include a third SO counterion. In our default setup, only one C-terminus of the Lys dimer interacts with a QM-level sulfate ion, while the second interfaces with a classical one. This modified setup allows us to evaluate whether explicitly treating both counterions at the QM level significantly alters the key vibrational modes involved in non-radiative decay. To this end, we selected five equilibrated initial conditions and re-equilibrated the system in the ground state for 1 ps using the enlarged QM region, following the same protocol as the second equilibration step described above. The final nuclear positions and velocities of these new QM/MM trajectories were then used to carry out 1 ps NAMD trajectories. All NAMD parameters were kept identical to those employed in the main simulations.
S3 Supplementary Text
S3.1 Determination of the Critical Aggregation Concentration
Lys samples for the experimental morphological and optical characterization were prepared by dissolving the amino acid powder in double distilled water at the concentration of 1 g/mL (6.8 mol/L), which corresponds to the maximum of its solubility. Under these conditions, an increase of the volume of was observed, and the resulting solution exhibited a slight yellow coloration. The pH of the resulting solution was basic (pH = 11.8). Given the known ability of single amino acids to self-assemble into supramolecular structure98, the formation of Lys aggregates was investigated using two environment-sensitive fluorescent probes: ANS and pyr, as shown in Fig. S199. Both dyes are known to be able to change their fluorescence properties as function of the surrounding chemical environment (hydrophobic or hydrophilic). ANS shows a characteristic emission peak between 460-480 nm in hydrophobic environments, such as the inner micellar compartment100 or aromatic peptide aggregate spine101. The fluorescence of pyr, on the other hand, is characterized by five vibrational bands, where the intensity ratio between the first (I1 at 373 nm) and third (I3 at 383 nm) bands serves as a sensitive indicator of environmental polarity.
The CAC of Lys was estimated from the inflection points of fluorescence titration curves (Fig. S1). CAC values determined using ANS (0.915 mol/L) and pyr (0.888 mol/L) fluorophores are in good agreement, supporting the reliability of the measurements. These values are slightly higher than the previously reported estimate of 0.5 mol/L by Stagi et al. 32, who identified a change in the fluorescence intensity slope as a function of concentration: at lower concentrations, intensity increased linearly, while at higher concentrations (from 0.4 mol/L), the rate of increase slowed, suggesting the formation of larger, less fluorescent aggregates. The fact that the ANS and pyr calculated values are below the Lys solubility (solubity/CAC 7.5) confirms that Lys aggregates, or smaller clusters, are present under the experimental conditions investigated.
S4 Supplementary Figures